Question 02: (20 pts) Two catalysts are being analyzed to determine how they affect the mean yield of a chemical process. Specifically catalyst 1 is currently in use, but catalyst 2 is acceptable. Since catalyst 2 is cheaper, it should be adopted, providing it does not change the process yield. A test is run in the pilot plant and results in the data shown in the table below Observation Number | Catalyst 1 (%) Catalyst 2 (%) 91.50 94.18 92.18 95.39 91.79 89.07 94.72 89.21 89.19 90.95 90.46 93.21 97.19 97.04 91.07 92.75 2 4 7 a) Is there any difference between the mean yields? Use a 0.05 and assume equal variances.(10 pts) se that if catalyst 2 produces a mean yield that differs from the mean yield of catalyst l by 40% we would like to reject the null hypothesis with probability at least 085. What sample size is required? (1o pt)
11. I have seen some Western parents teach their children to be independent at an early age. This explains why Western children are very independent. Which fallacy does this argument contain? A. Red Herring B. Inappropriate appeal to authority C. Hasty generalization D. Questionable cause
Answer #1

At first put all the data in an Excel spreadsheet and calculate the mean and standard deviation using the AVERAGEO, STDEV.S()

The sample standard deviation of catalyst 2(%): |(89.19-92.73) +(90.95-92.73)? +...+(92.75-92.73)2 8-1 12.55+3.18 +...+0.0003

The degrees of freedom are: = n +12 - 2 = 8+8-2 = 14 At 14 degrees of freedom and 0.05 level of significance, the critical va